# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v2

@@ 3.1. Material: The Narrative Framework
For our study, we selected the Harry Potter world as the narrative framework for our FI method. The Harry Potter series follows the journey of an orphaned boy who discovers he is a wizard and, as a result, is invited to attend “Hogwarts School of Witchcraft and Wizardry,” where he forms deep friendships and gradually uncovers his destiny as he clashes with the story’s antagonist.
Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI interviews. The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement. The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells. Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies. Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building. The characters’ experiences or life transitions, which parallel real-world moves between schools or locations, allow for exploring how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
@@ 3.2. Procedure
The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI method. The interviews averaged around 90 minutes. We began the interview by asking which 3D games they had played and what they liked about those experiences. We also asked questions about Harry Potter, such as their favorite book in the series. This served to break the ice and help them start immersing themselves in the Harry Potter world.
Following this introduction, we presented the FI narrative: participants were asked to envision ways to support social connection within the Hogwarts world as a non-Wizard (i.e., “Muggle” (mug, [n. d.])) trying to stay connected with everyone across Hogwarts and the Muggle world. We then shared a Miro (mir, [n. d.]) board (Figure 1) where participants could sort their friends and family into categories inspired by the Harry Potter series, such as fictional houses (e.g., Gryffindor, Hufflepuff), teams (e.g., Quiddich Team members), and old friends (e.g., muggles), reflecting fictional houses, teams, and communities from the story. This activity encouraged them to think of their relationships within the Harry Potter context and was intended to further enhance immersion in the fictional design space. Next, we asked general questions about their experiences with social media, focusing on what they found meaningful or meaningless and their vision for an ideal social media platform.
During the design phase, participants were prompted to describe the capabilities their ideal magical social media would have and what aspects of their lives they would choose to share. Throughout this phase, we sought to keep participants immersed in the narrative, referring to existing platforms as “Muggle social media” and maintaining consistent engagement with the Hogwarts setting, as outlined by the fictional inquiry guidelines. Finally, we concluded by asking participants to reflect on how the social media they envisioned for Hogwarts compared to Muggle social media. The semi-structured interview protocol is available in Appendix A.
@@ 3.3. Participants and Recruitment
We used a combination of purposive and convenience sampling to recruit participants for our study. The first and second authors shared the recruitment materials (that included a link to a screener survey) on their social media accounts, while the first author also distributed flyers across our institution’s campus.
Participants were required to be between 15 and 24 years old, in accordance with UN (United Nations, [n. d.]) and WHO (who, [n. d.]) definitions for youth, and to consent to be video or audio recorded during a 90-minute interview. Video recording was requested following the FI method guidelines, enabling us to analyze gestures if necessary. We included both teenage and young adult participants (ages 15 to 24) to capture perspectives from distinct life stages, as these groups often navigate significant transitions and share a focus on identity exploration, self-presentation, and relationships.
They also had to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox). While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
@@
A total of 84 individuals completed the screener survey. We reached out to the 48 respondents who met the recruitment criteria, and 23 participants provided consent and completed the interview procedure. The mean age of the participant group was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) identifying as women, 3 (13.0%) as men, 2 (8.7%) as non-binary, and 3 (13.0%) that were gender fluid. The full list of participant information is available in Table 1.
@@ 3.4. Data Analysis
We analyzed the interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides a systematic yet adaptable framework for interpreting qualitative data. The method’s independence from pre-existing theoretical frameworks made it particularly suitable for our study, given our study’s exploratory focus on youth-envisioned social media design through the FI method. It enabled us to effectively capture and analyze the diverse perspectives shared by our participants.
Our analysis began with the first two authors independently coding the same two transcripts line by line. During this phase, the codes were descriptive and closely adhered to the data. The authors convened to discuss and resolve any discrepancies in their coding. Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024). After four coding iterations, we reached saturation in our codes. Using this final set of codes, the first author proceeded to code all the transcripts. All versions of the codes and the final codebook are available in the supplementary materials.
@@ 3.5. Ethical Considerations
Given J.K. Rowling’s controversial statements regarding transgender individuals (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in our recruitment material and screener survey and reiterated our position during the interview procedure.
@@ 4. Results
Informed by their experiences with 3D games and familiar media worlds, our participants imagined forms of social interaction that extend beyond what is typically supported on mainstream platforms like Instagram. Through the speculative frame of Social Media at Hogwarts (SMH), they explored how digital environments might enable more embodied, relational, and emotionally meaningful forms of connection. Their visions converged around six interrelated themes, which, collectively, contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs, ultimately expanding their perception of the horizons for social connection on such platforms.
@@ 4.1. Intuitive Social Navigation
Youth described seeking interactions that feel natural, embodied, and socially meaningful, reflecting the way people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape how they connect with others, express themselves, and manage privacy in intuitive and socially familiar ways.
@@ 4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Participants described how embodied, immersive qualities such as realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space that fosters presence, emotional connection, and socially meaningful interaction. In contrast, current social media (which we will refer to as “Muggle Social Media” or MSM for short), was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 similarly noted that MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” P03 described SMH as “more immersive and more awesome” than MSM, which they saw as inherently limited “because it’s 2D.” P16 emphasized that meaningful social connection requires “talk[ing] to you and see[ing] your face,” something they believed SMH could make possible.
For others, the fulfillment came from the embodied experience of physical actions. P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They also added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 similarly reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” As P06 put it, SMH “feels more real” because you “walk around a physical world” and “feel like you’re actually in the area of these people.”
@@
For P01, 3D interaction not only supported more embodied engagement but also introduced a felt presence that lingered, even when no direct interaction took place.
@@ 4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed that using spatial and visual cues modeled after real-world environments could make managing access and privacy feel more intuitive, socially natural, and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
@@
While managing visibility across multiple rooms might seem more complex, participants believed this approach would feel intuitive and natural. P11 suggested that spatial contexts could lower the efforts required for privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
@@ 4.2. Shared Experiences
Participants emphasized that shared experiences, whether through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
@@ 4.2.1. Fostering Shared Memories Through Collaborative Activities
Participants described shared activities as a meaningful foundation for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. For example, while playing a simple emoji game was described as “just something fun we can both do,” it lacked a true shared experience since “we don’t actually see each other’s screens” and would just “screenshot our scores and send them to each other.” By contrast, they described Roblox as a space that fosters shared moments that become lasting memories, like “do you remember that time when that kid in the voice chat said this,” or “when we got scared by that jump scare.” They explained, “it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
Similarly, P12 described the value of collaborative projects, such as “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting it with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than simply “posting” content individually.
The value of shared experiences and memories was not limited to completing tasks or reaching goals. Even simple activities like walking together were seen as meaningful opportunities for connection. P03 imagined moments where you could just “walk with your friend,” as if taking a “nature walk” (P14). P21 echoed this, describing how walking itself could spark memorable conversations and shared moments: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
@@ 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Participants also saw low-pressure, ambient forms of social interaction, such as co-existing in shared spaces or listening to music together, as natural and organic ways to build connections and shared memories. P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background. They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations. P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.” Even “mindless scrolling” became meaningful when shared. P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music would often facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They came up with the idea reflecting on their own experience:
@@
They added that music could also serve as a “conversation starter” while also providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
@@ 4.3. Communal Ambience
SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships. Participants described several design features, such as neighborhood structures, shared gathering spaces, and stronger boundary controls, that could help cultivate this kind of communal, low-pressure environment.
@@ 4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to naturally foster social interaction and relationship-building. P22 described in detail:
@@
Youth envisioned SMH as also reflecting the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties. Specifically, some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction. One participant elaborated:
@@
Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them. This design approach of aligning physical landscapes with emotional closeness was seen as a way to streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”

### v3

@@ 3.2. Material: The Narrative Framework
We selected the Harry Potter world as our FI narrative framework. The series follows an orphaned boy who discovers he is a wizard and attends “Hogwarts School of Witchcraft and Wizardry,” forming deep friendships while uncovering his destiny and clashing with the antagonist.
Among numerous magical worlds, Harry Potter proved particularly suitable for FI interviews. Its extensive popularity and cultural presence provide shared understanding, facilitating recruitment and engagement. By 2023, 98% of US youth aged 18-27 identified as casual or avid fans, with half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). The book series sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and was translated into over 80 languages by 2020 (The Rowling Library, 2020).
Set in modern England with magical elements like enchanted objects and spells, this context allows incorporating or discarding aspects of contemporary life, including social technology designs and constraints. Young protagonists and coming-of-age narrative align with our focus on youth identity and relationship building. Character transitions parallel real-world moves between schools or locations, enabling exploration of how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging suits examining youth identity expression and self-presentation in virtual spaces.
@@ 3.3. Procedure
The first author conducted all interviews over Zoom, with the second author joining one session to validate the FI method. Interviews averaged 90 minutes. We began by asking about 3D gaming experiences and Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (mug, [n. d.])) connecting across Hogwarts and Muggle worlds. We shared a Miro (mir, [n. d.]) board (Figure 1) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles). This encouraged thinking about relationships within the Harry Potter context and enhanced fictional design space immersion. We then asked about social media experiences, focusing on meaningful versus meaningless aspects and ideal platform visions.
During the design phase, participants described their ideal magical social media capabilities and what life aspects they would share. We maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines. On the other hand, we encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?” and “If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)” We concluded by asking participants to compare their envisioned Hogwarts social media with Muggle social media. The semi-structured protocol is in Appendix A.
@@ 3.4. Participants and Recruitment
We used purposive and convenience sampling for recruitment. The first and second authors shared recruitment materials (including screener survey links) on social media, while the first author distributed campus flyers.
Participants needed to be 15-24 years old per UN (United Nations, [n. d.]) and WHO (who, [n. d.]) youth definitions and consent to video/audio recording during 90-minute interviews. Video recording followed FI guidelines, enabling gesture analysis if needed. We included teenagers and young adults to capture perspectives from distinct life stages sharing focus on identity exploration, self-presentation, and relationships.
Participants needed sufficient Harry Potter familiarity (agreeing with “I remember some details about the plot, characters, and magical aspects”) and prior 3D gaming experience (e.g., Minecraft, Roblox). We did not require extensive knowledge of Harry Potter because we felt that would be an unnecessary bias that is not particularly helpful in envisioning new social media design. While this made it somewhat difficult for a majority of participants to remember specific artifacts from the Harry Potter world, they knew enough to recognize the complete and extensive freeing extent of the magical powers in that setting. Though we imposed no social media experience constraints, all participants had some social media engagement.
@@
Eighty-four individuals completed the screener survey. We contacted 48 who met criteria; 23 provided consent and completed interviews. Mean age was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) women, 3 (13.0%) men, 2 (8.7%) non-binary, and 3 (13.0%) gender fluid participants. While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel, 2019), we note that our participant demographics skew toward women. Full demographic information appears in Table 1.
(Lenhart, 2015b, c)
@@ 3.5. Data Analysis
We analyzed interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides systematic yet adaptable qualitative data interpretation. Its independence from pre-existing theoretical frameworks suited our exploratory focus on youth-envisioned social media design through FI, enabling effective capture and analysis of diverse participant perspectives.
Since not all participants agreed to have their videos on during interviews, we did not conduct video analysis of gestures that participants sometimes used when describing their ideas. Additionally, drawing was not part of our protocol because we have experienced in prior interviews for other studies that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination. All data therefore consisted of transcripts from verbal interviews.
Analysis began with the first two authors independently coding the same two transcripts line-by-line using descriptive codes closely adhering to data. Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atl, 2024). After four iterations, we reached code saturation. The first author then coded all transcripts using the final code set. All code versions and the final codebook are in supplementary materials.
@@ 3.6. Ethical Considerations
Given J.K. Rowling’s controversial transgender statements (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in recruitment materials and surveys, reiterating our position during interviews.
@@ 4. Results
Drawing from their experiences with 3D games and familiar media worlds, participants imagined social interactions extending beyond mainstream platforms like Instagram. Through the speculative frame of exploring a “new magical ‘thing’ for social communication” for Hogwarts students, they envisioned how digital environments might enable more embodied, relational, and emotionally meaningful connection. Their visions converged around six interrelated themes that collectively contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs. We refer to this synthesized concept as “Social Media at Hogwarts,” or SMH, throughout this paper.
@@ 4.1. Intuitive Social Navigation
Youth sought interactions that feel natural, embodied, and socially meaningful, reflecting how people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy management in intuitive, socially familiar ways.
@@ 4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Participants described how embodied, immersive qualities like realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space fostering presence, emotional connection, and meaningful interaction. In contrast, current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation, explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
For others, fulfillment came from embodied physical actions. P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing,” explaining “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
@@
For P01, 3D interaction supported more embodied engagement and introduced a felt presence that lingered, even without direct interaction.
@@ 4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed using spatial and visual cues modeled after real-world environments to make managing access and privacy feel more intuitive, socially natural, and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
@@
While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive and natural. P11 suggested spatial contexts could reduce privacy control efforts, stating “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
@@ 4.2. Shared Experiences
Participants emphasized that shared experiences, through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
@@ 4.2.1. Fostering Shared Memories Through Collaborative Activities
Participants described shared activities as meaningful foundations for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” they described Roblox as fostering lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
The value of shared experiences extended beyond completing tasks. P21 described how walking could spark memorable conversations: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
@@ 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Participants saw low-pressure, ambient social interaction, like co-existing in shared spaces or listening to music together, as natural ways to build connections and shared memories. P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual study room with Lo-Fi music. These spaces would mirror Discord servers’ 24/7 activity, offering comforting presence even without direct conversation. Even “mindless scrolling” became meaningful when shared, as P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music often facilitated these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs,” reflecting on their experience:
@@
They added that music could serve as a “conversation starter” while providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
@@ 4.3. Communal Ambience
SMH was envisioned as a platform for intentional, engaging interactions with close friends, making it naturally less prone to social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features fostering organic, intentional, personalized interactions and closer friendships. They described several design features–neighborhood structures, shared gathering spaces, and stronger boundary controls–that could cultivate this communal, low-pressure environment.
@@ 4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
Participants wanted SMH’s 3D world to mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster social interaction and relationship-building. P22 described in detail:
@@
Youth envisioned SMH reflecting real-world relationship dynamics, using physical distances within the landscape to represent social tie strength. Some participants envisioned SMH resembling a college campus or neighborhood, where physical closeness naturally fosters interaction. One participant elaborated:
@@
P01 proposed a tiered layout where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be farther away, requiring a “little hike” to reach. This approach of aligning physical landscapes with emotional closeness would streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”